A binomial distribution consists of 120 trials, with a success probability of 43% . determine the mean and standard deviation


1
Expert's answer
2021-05-21T12:07:01-0400

mean "\mu = np"


"\mu = 120 \times 0.43 = 51.6"


standard deviation= "\sigma"


"\sigma^2= np(1-p)= 120\times 0.43(1-0.43)"


"\sigma^2= 29.412"


"\sigma" = "\sqrt{29.412}= 5.423"